OpenStudy (kirbykirby):

Is this valid to write to say that x belongs to the negative integers, and the element zero? \(x \in \mathbb{Z}^- \cup \left\{ {0}\right\}\) ?

OpenStudy (anonymous):

Yes, as long as \(\mathbb{Z}^-=\text{the set of all negative integers}\). You might see the same set written as \(\mathbb{Z}_-\), or possibly \[\left\{x\in\mathbb{Z}:x\le0\right\}\]

OpenStudy (kirbykirby):

ok thank you =]